OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = −x + 7 10 5 over 2 2 over 5 20

OpenStudy (anonymous):

This the begining equation correct? \[\frac{ -1 }{ 2x+2 } = -x+7\]

OpenStudy (anonymous):

or is it (1/2)x + 2 = -x + 7 ?

OpenStudy (anonymous):

Didn't you just post this and get 10?

OpenStudy (anonymous):

-1/2x+2=-x+7

OpenStudy (anonymous):

multiply both sides by -2 to get rid of the (-1/2)x

OpenStudy (anonymous):

i didnt understand it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[-\frac{1}{2}x+2=-x+7\] \[x-4=2x-14\] \[x-4+(14)=2x-14+(14)\] \[x-10=2x\] \[-10=-x\] \[?=x\]

OpenStudy (anonymous):

so ten

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Need help with anything else?

OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.